OpenStudy (swaggyswags):

What is the equation, in slope-intercept form, of the line that passes through (0, 6) and has a slope of 4? y = 4x − 6 y = 4x + 6 y = 6x − 4 y = 6x + 4

OpenStudy (will.h):

you do know that the slope is 4 so it is either option A or B

OpenStudy (will.h):

you know that or no?

OpenStudy (swaggyswags):

Yeah it's got a slope of 4

OpenStudy (swaggyswags):

It'd be B right? Because it's a positive slope?

OpenStudy (will.h):

okay do you know this formula y - y1 = m(x - x1) we have coordinate of (0,6) and slope of 4 so therefor y - 6 = 4(x - 0) y - 6 = 4x - 0 y = 4x + 6 Hope that helps

OpenStudy (swaggyswags):

Thanks man :) Could you help me with a few more?

OpenStudy (will.h):

tag me on new post

OpenStudy (swaggyswags):

Alright